OpenStudy (anonymous):

If m DFC = 40° and m arc CD= 55°, then m arc bg =

OpenStudy (anonymous):

heya this one seems very similar to the last one is it the same formula or a different one?

OpenStudy (anonymous):

https://media.glynlyon.com/g_geo_2012/7/group177.gif

OpenStudy (anonymous):

oh there's the link btw

ganeshie8 (ganeshie8):

this is lil tricky

ganeshie8 (ganeshie8):

lets work it, if, m DFC = 40 m BFC = ?

OpenStudy (anonymous):

not sure

ganeshie8 (ganeshie8):

open the pic i attached

ganeshie8 (ganeshie8):

that green angle is m DFC = 40, right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

m BFC = 180-40 cuz, BD is a straight line !

OpenStudy (anonymous):

mmmk so 140

ganeshie8 (ganeshie8):

before we move further solving the problem, pls make sure you understand how we got m BFC = 180-40 = 140